OpenStudy (anonymous):

What is the volume of a sphere with a radius of 6 inches? 112Ï€ cubic inches 288Ï€ cubic inches 290Ï€ cubic inches 300Ï€ cubic inches

jimthompson5910 (jim_thompson5910):

Volume of a Sphere \[\Large V = \frac{4}{3}\pi*r^3\] plug in r = 6 and simplify

OpenStudy (anonymous):

@jim_thompson5910 i got a answer i just wanna make sure other people get the same outcome i did.

jimthompson5910 (jim_thompson5910):

what is \[\Large \frac{4}{3}*r^3\] equal to when r = 6

OpenStudy (anonymous):

i got 216 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

close, but incorrect

OpenStudy (zale101):

r^3=216 but you didn't finish solving the equation.

OpenStudy (zale101):

6^3=216

OpenStudy (anonymous):

\[288 \pi \]

OpenStudy (unklerhaukus):

\[V_\text{sphere}(r)=\tfrac43\pi r^3\] \[V_\text{sphere}(6\,[\text{in}])\\ =\tfrac43\pi(6\,[\text{in}])^3\\ =\tfrac43\pi\cdot 216\,[\text{in}]^3\\ =(\tfrac43\times216) \pi\,[\text{in}]^3\\=\]